OpenStudy (anonymous):

The first four terms of a sequence are shown below. 8, 5, 2, -1 Which of the following functions best defines this sequence? f(1) = 8, f(n + 1) = f(n) + 3; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) + 5; for n ≥ 1 f(1) = 8, f(n + 1) = f(n) - 3; for n ≥ 1

OpenStudy (anonymous):

You are subtracting by three in the sequence. There is only one option that will let you subtract by 3. That is the last option, D

OpenStudy (anonymous):

so its d

OpenStudy (shamil98):

he just said that.

OpenStudy (anonymous):

thanks if you don't mind i have more questions can u help me with them

OpenStudy (anonymous):

Yeah sure.

OpenStudy (anonymous):

Which of the following statements best describes the effect of replacing the graph of y = f(x) with the graph of y = f(x) + 20? The graph of y = f(x) will shift left 20 units. The graph of y = f(x) will shift down 20 units. The graph of y = f(x) will shift up 20 units. The graph of y = f(x) will shift right 20 units.

OpenStudy (anonymous):

Well you know that + 20 is the y intercept right?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

That leaves with you shifting up or down. Because it is positive, which way do you think you are going to shift?

OpenStudy (anonymous):

up

OpenStudy (anonymous):

thanks hers another

OpenStudy (anonymous):

Every week Annie plants 2 saplings in her garden. The function below shows the total number of saplings f(w) in her garden after w weeks. f(w) = 2w + 3 Which statement best describes the y-intercept of the graph of the function? It is equal to 3 and represents the original number of saplings in Annie's garden. It is equal to 2 and represents the original number of saplings in Annie's garden. It is equal to 3 and represents the number of saplings in Annie's garden at the end of the first week. It is equal to 5 and represents the number of saplings in Annie's garden at the end of the first week.

OpenStudy (anonymous):

this one I put the 1st one

OpenStudy (anonymous):

Correct

OpenStudy (anonymous):

Choose the graph below that correctly represents the equation 2x + 4y = 24.

OpenStudy (the_fizicx99):

Put it into y =mx + b then plot b ( y intercept_ and do rise/run

OpenStudy (the_fizicx99):

Subtract x from both sides then divide everything by 4 to just have y

OpenStudy (anonymous):

so its A

OpenStudy (the_fizicx99):

How did you get A

OpenStudy (anonymous):

subtracted 2 on both sides and it became 4y=22 then dived and got 5.5

OpenStudy (the_fizicx99):

No... 2x + 4y = 24 -2x -2x 4y = -2x + 24 /4 /4 /4 y = -1/2 + 6

OpenStudy (the_fizicx99):

PLot that

OpenStudy (the_fizicx99):

Plotting it at +6 on the y, then you would go down 1 and right 2

OpenStudy (the_fizicx99):

So yes it A

OpenStudy (anonymous):

1 more question Choose the slope-intercept equation of the line that passes through the point shown and is perpendicular to the line shown.